# zkFuzz: Foundation and Framework for Effective Fuzzing of Zero-Knowledge Circuits

### v1

@@ 4.2. Guided Search
zkFuzz uses the smoothed error function transformed from the constraints $\mathcal{C}$ as the fitness score to find an execution trace that satisfies the constraints while producing a different output from the original. This means that we guide the mutation towards under-constrained circuits, and we adopt this design since under-constrained circuits tend to cause more critical outcomes than over-constrained circuits (Chen et al., 2024).
The constraints $\mathcal{C}$ usually consist of a set of polynomial equalities, $\bigwedge_{i=1}^{|\mathcal{C}|}(a_{i}==b_{i})$, where $a_{i}$ and $b_{i}$ are quadratic expressions, and $|\mathcal{C}|$ denotes the number of those equations. Following (Chen et al., 2022b), we convert each equality into an error term $|a_{i}-b_{i}|$, which evaluates to 0 when the constraint is satisfied and a positive value when violated. The total error for a given trace is defined as the sum of these errors if the trace produces a different output from the original. Otherwise, the error is set to $\infty$, reflecting the failure to find a diverging execution. The final fitness score of a mutated program is determined by evaluating its error over multiple inputs and selecting the minimum score. Our min-sum method differs from prior max-only method (Chen et al., 2022b) because zkFuzz jointly mutates both programs and inputs.
In each iteration, zkFuzz uses roulette-wheel selection (Razali et al., 2011), a widely adopted method in genetic algorithms, to generate new program mutants or perform crossover between mutants from the previous iteration. The selection is biased toward mutants with lower scores, as lower values indicate higher fitness. zkFuzz represents a program mutant as a map from candidate mutation locations to the actual mutations applied, so crossover is implemented as a straightforward merging of the maps from the two parent mutants.
@@
Target Selectors. We further improve the search of zkFuzz by incorporating target selectors, which guide both input generation and program mutation. Unlike traditional targeted fuzzers—which primarily detect issues like memory access violations by finding inputs that reach specific code blocks (Weissberg et al., 2024)—our target selectors are designed to maximize the odds of uncovering discrepancies between the computation program $\mathcal{P}$ and the circuit constraints $\mathcal{C}$.
@@
One of the key target selectors is skewed distribution, which generates input data randomly while biasing the distribution to regions more likely to contain vulnerabilities. Specifically, empirical analysis reveals that many bugs stem from edge-case inputs, particularly values such as 00 and the prime modulus $q$. To effectively target these scenarios, our tool samples inputs and substitutes constants in unconstrained assignments using a skewed distribution that increases the likelihood of uncovering hidden vulnerabilities. Specifically, values are sampled from predefined ranges with the following probabilities: binary values (00 and $1$) with 15%, small positive integers ($2$ to $10$) with 34%, large values near the field order ($q-100$ to $q-1$) with 50%, and all other values ($11$ to $q-101$) with 1%. By biasing input selection toward these critical edge cases, our approach enhances the detection of subtle bugs that might otherwise go unnoticed. The impact of the choice of skewed distribution can be found in § 6.4.
@@
In addition, our zkFuzz leverages static analysis to mutate computation programs and generate inputs more likely to trigger unexpected behaviors. For example, one of the main root causes of real-world under-constrained circuits is the zero-division pattern, wherein the division’s numerator and denominator may be zero (Wen et al., 2024). In this scenario, a naive constraint for computing $y=x_{1}/x_{2}$ might be written as $y\cdot x_{2}=x_{1}$. However, if both $x_{1}$ and $x_{2}$ are zero, any value of $y$ will satisfy this constraint, leaving the circuit under-constrained. Therefore, when this pattern is detected, increasing the chance that generated values for $x_{1}$ and $x_{2}$ are zero can help reveal these vulnerabilities. Details of these static-analysis-based target selectors are provided in § 5.2.
5. zkFuzz Implementation
@@
We implemented zkFuzz for Circom in about 4,500 lines of Rust. It begins with an enhanced parser that decomposes Circom programs into an Abstract Syntax Tree (AST) for the computation $\mathcal{P}$ and corresponding circuit constraints $\mathcal{C}$. A mutator generates program variants ($\mathcal{P}^{\prime}$), and an emulator produces execution traces for these variants using artificial inputs. A bug detector checks if these traces satisfy the original circuit constraints based on TCCT. Target selectors strategically guide input generation and program mutation towards likely bugs. Finally, mutants’ fitness scores drive selection and crossover in subsequent generations (Fig. 2).
5.1. Circom Program Mutation
@@
zkFuzz mutates Circom programs using the following methods.
@@
(b) Weak-Assignment Transformations: We randomly alter the right-hand side of weak assignment operators (<--). Recall that <-- does not introduce constraints, meaning that modifying its right-hand side does not affect the constraints derived from the original Circom file. Based on the popular mutation strategies used in mutation testing (Papadakis et al., 2019), we apply two types of mutations: 1) By default, we replace the right-hand expression with a random value over the finite field sampled from the skewed distribution (see § 4.2); 2) If the right-hand side contains an operator, we may also substitute it with a different operator with the same category, where we categorize the operators into arithmetic, bitwise, and logical operators. Note that mutating the strong assignment (<==) is ineffective, as it is likely to violate the associated condition.
We implement two optional mutation methods, random value addition and expression deletion. We show detailed results in Appendix D.2.
5.2. Target Selection with Static Analyzers
Finding bugs in ZK circuits is particularly challenging due to the vast search space imposed by large finite fields. To overcome this challenge, zkFuzz leverages static analysis for selecting promising targets that guide both program mutation and input generation.
Table 3. Summary of static target selectors used in zkFuzz.
@@ 5.2. Target Selection with Static Analyzers
Hash-Check Input Generation h(x1) $=$ x2
White List Prog. Mutation IsZero, Num2Bits
Zero-Division Pattern. As discussed in § 4.2, the zero-division pattern is a common source of under-constrained circuits where both the division’s numerator and denominator can be zero. Since Circom lacks native support for non-quadratic constraints like division, programmers must manually specify equivalent quadratic constraints. Notably, when the denominator is zero, Circom treats the division as yielding zero.
Code 6 illustrates a simple case of this issue: when both numerator and denominator are zero, the multiplication constraint (see line 7) is trivially satisfied for any value of y, due to missing checks for zero-division. Appendix B also presents a real-world circuit example from circomlib, where zkFuzz successfully detects an under-constrained condition using only the zero-division selector.
@@
(Degree 2) Consider the quadratic expression $ax^{2}+bx+c\equiv 0\bmod{q}$. A solution for $x$ is given by $x\equiv\frac{-b+\sqrt{b^{2}-4ac}}{2a}\bmod{q}$, provided that a square root of $b^{2}-4ac$ exists modulo $q$. Its existence can be verified using Euler’s criterion, and the Tonelli-Shanks algorithm (Tonelli, 1891; Shanks, 1972) is then employed to compute $\sqrt{b^{2}-4ac}$, yielding the solution.
@@
Hash-Check Pattern. A common pattern in ZK circuit is the hash-check pattern, where a circuit verifies that the hash of some input data matches a provided hash value (see Code 7). Naive fuzzing with program mutation may struggle to generate valid input-hash pairs, especially when cryptographic hash functions are involved. To overcome this limitation, we employ a heuristic: with a certain probability if an equality constraint (e.g., A == B) is encountered where A is an input signal that has not been assigned at this point, we update the assignment of A to match the value of B. This adjustment increases the likelihood that the input data will satisfy the hash-check constraint. A real-world example is shown in Appendix B, where zkFuzz identifies the circuit as under-constrained using only the Hash-Check selector.
@@
2 signal input x1;
3 signal input x2;
@@
2 signal input x1;
3 signal input x2;
4
@@
White List. Existing research and public audit reports have formally proven the safety of certain commonly used circuits, even if they contain weak assignments. Since mutating these circuits would immediately lead to constraint violations, zkFuzz allows specifying a white list of circuit templates to be skipped during the mutation phase, concentrating the mutation efforts on less well-understood or more error-prone components. In this study, we include the two most commonly used circomlib circuits with weak assignments, IsZero and Num2Bits, in the white list.
Collectively, these target selectors enable zkFuzz to navigate the enormous search space inherent in ZK circuits efficiently. By strategically guiding both program mutation and input generation, they significantly enhance our ability to detect critical vulnerabilities.
6. Evaluation
In this section, we present a comprehensive experimental evaluation designed to address the following research questions:
@@ 6. Evaluation
  • RQ4: What is the individual contribution of each heuristic and static analyzer in zkFuzz to efficient bug detection?
@@
We further consider the following questions in Appendix D:
  • RQ6: Is TCCT more general than existing formulations?
@@
  • RQ9: How many over-constrained circuits are observed when the constraint_assert_disabled flag is set?
@@
Benchmark Datasets. We evaluate zkFuzz on 354 real-world test suites written in Circom. Our benchmark extends the dataset from ZKAP (Wen et al., 2024) by adding new test cases from 30 additional projects. Following Pailoor et al. (Pailoor et al., 2023), we classify the circuits into four categories based on the number of clauses in the constraint, denoted as $|\mathcal{C}|$ (e.g., for $\mathcal{C}(x,z,y):=(x*x=z)\land(z*z=y)$, $|\mathcal{C}|=2$). The categories are defined as follows: Small ($|\mathcal{C}|<100$); Medium ($1000100\leq|\mathcal{C}|<$ 1000); Large ($100001000\leq|\mathcal{C}|<$ 10000); Very Large: ($10000\leq|\mathcal{C}|$). Detailed benchmark stats are in Appendix C.
@@
Configurations. We set a timeout of 2 hour and a maximum of 50000 generations for all our tests. We also set the number of program mutants per generation to 30, each paired with generated inputs of the same size. Program mutation and crossover operations are applied with probabilities of 0.3 and 0.5, respectively. By default, the right-hand side of a weak assignment is replaced with a random constant. However, if it consists of an operator expression, the operator is randomly substituted with a probability of 0.1. When a zero-division pattern is detected, we analytically solve for the variable and substitute it with the computed solution with a probability of 0.2. Those parameters are determined with reference to prior works on genetic algorithm mutation testing (Shin et al., 2019; Hassanat et al., 2019; Lin et al., 2003), and we empirically demonstrate the robustness of zkFuzz across different hyperparameter settings in § 6.5. We run zkFuzz with five different random seeds. All experiments were run on an Intel(R) Xeon(R) CPU @ 2.20GHz and 31GB of RAM, running Ubuntu 22.04.3 LTS.
@@
Baseline Approaches. We compare zkFuzz against four state-of-the-art Circom bug detection tools: Circomspect (of Bits, 2024), ZKAP (Wen et al., 2024), Picus (Pailoor et al., 2023), and ConsCS (Jiang et al., 2025). We use two SMT solvers, Z3 (De Moura and Bjørner, 2008) and CVC5 (Barbosa et al., 2022) for Picus. To ensure a fair comparison, we do not count warnings related to IsZero and Num2Bits as false positives, since zkFuzz explicitly whitelists these templates. Additionally, ZKAP’s unconstrained signal (US) detector flags any circuit containing unused signals as unsafe. Although the original ZKAP paper considers unused inputs problematic, our work does not regard them as bugs (see § 7). Moreover, as noted in (Wen et al., 2024), the primary source of ZKAP’s false positives is its unconstrained sub-circuit output (USCO) check, which targets unused component outputs. Therefore, we exclude both US and USCO from our main results to improve ZKAP’s precision. Note that the original Circomspect does not analyze internally called templates recursively. Lastly, zkFuzz-Naive is a simplified variant of zkFuzz that mutates both strong and weak assignments, uses a constant function as the fitness function, and disables all static-based target selectors listed in Tab. 3.
@@
Table 4. The number of unique bugs detected by each tool, categorized by circuit size (measured by the size of constraint, $|\mathcal{C}|$). TP and FP denote true positive and false positive, respectively. Precision is computed as TP / (TP + FP). Max / Min refers to the highest and lowest number of detected bugs across five different random seeds. The results show that zkFuzz significantly outperforms the existing methods. The magnitude of TP and Precision, and FP are highlighted in blue and red, respectively.
Constraint
Size
#Bench #Bug Circomspect (of Bits, 2024) ZKAP (Wen et al., 2024)
Picus (Pailoor et al., 2023)
(Z3 / CVC5)
ConsCS (Jiang et al., 2025)
zkFuzz (Naive)
(Max / Min)
zkFuzz (Ours)
(Max / Min)
TP FP TP FP TP FP TP FP TP FP TP FP
Small 144 41    37 16 32 9 19 / 20 0   14 0 40 / 39 0 41 / 41 0
Medium 51 6 3 6 0 12 0 0 0 0 6 / 6 0 6 / 6 0
Large 76 16 6 12 3 2 0 0 0 0 12 / 12 0 15 / 15 0
Very Large 83 6 5 18 2 2 0 0 0 0 2 / 2 0 4 / 3 0
Total 354 69 51 52 37 25 19 / 20 0 14 0 60 / 59 0 66 / 65 0
Precision 0.50 0.60 1.00 1.00 1.00 1.00
6.1. RQ1: Effectiveness of zkFuzz

### v2

@@ 4.2 Guided Search
zkFuzz uses the smoothed error function transformed from the constraints $\mathcal{C}$ as the fitness score to find an execution trace that satisfies the constraints while producing a different output from the original. This means that it guides the mutation towards under-constrained circuits, and we adopt this design since under-constrained circuits tend to cause more critical outcomes than over-constrained circuits [24].
The constraints $\mathcal{C}$ usually consist of a set of polynomial equalities, $\bigwedge_{i=1}^{|\mathcal{C}|}(a_{i}==b_{i})$, where $a_{i}$ and $b_{i}$ are quadratic expressions, and $|\mathcal{C}|$ denotes the number of those equations. Following [30], we convert each equality into an error term $|a_{i}-b_{i}|$, which evaluates to 0 when the constraint is satisfied and a positive value when violated. The total error for a given trace is defined as the sum of these errors if the trace produces a different output from the original. Otherwise, the error is set to $\infty$, reflecting the failure to find a diverging execution. The final fitness score of a mutated program is determined by evaluating its error over multiple inputs and selecting the minimum score. Our min-sum method differs from prior max-only method [30] because zkFuzz jointly mutates both programs and inputs.
In each iteration, zkFuzz uses roulette-wheel selection [46], a widely adopted method in genetic algorithms, to generate new program mutants or perform crossover between mutants from the previous iteration. The selection is biased toward mutants with lower scores, as lower values indicate higher fitness. zkFuzz represents a program mutant as a map from candidate locations to the actual mutations applied, so crossover is implemented as a straightforward merging of the maps from the two parent mutants.
@@
Target Selectors. We further improve the search of zkFuzz by incorporating target selectors, which guide both input generation and program mutation. Unlike traditional targeted fuzzers—which primarily detect issues like memory access violations by finding inputs that reach specific code blocks [47]—our target selectors are designed to maximize the odds of uncovering discrepancies between the computation program $\mathcal{P}$ and the circuit constraints $\mathcal{C}$.
@@
The default target selector of zkFuzz is skewed distribution, which generates input data randomly while biasing the distribution to regions more likely to contain bugs. Specifically, empirical analysis reveals that many bugs stem from edge-case inputs, particularly values such as 0 and the prime modulus $q$. To effectively target these cases, zkFuzz samples inputs and substitutes constants in unconstrained assignments using a skewed distribution that increases the likelihood of uncovering hidden vulnerabilities. Specifically, values are sampled from predefined ranges with the following probabilities: binary values (0 and $1$) with 15%, small positive integers ($2$ to $10$) with 34%, large values near the field order (q100q-100 to q1q-1) with 50%, and all other values with 1%. By biasing input selection toward these critical edge cases, our approach enhances the detection of subtle bugs that might otherwise go unnoticed. The impact of the choice of skewed distribution can be found in § 6.4.
@@
In addition, our zkFuzz optionally leverages static analysis to mutate computation programs and generate inputs more likely to trigger unexpected behaviors. For example, one of the main root causes of real-world under-constrained circuits is the zero-division pattern, wherein the division’s numerator and denominator may be zero [13]. In this scenario, a naive constraint for computing $y=x_{1}/x_{2}$ might be written as $y\cdot x_{2}=x_{1}$. However, if both $x_{1}$ and $x_{2}$ are zero, any value of $y$ will satisfy this constraint, leaving the circuit under-constrained. Therefore, when this pattern is detected, increasing the chance that generated values for $x_{1}$ and $x_{2}$ are zero can help reveal these vulnerabilities. Details of these static-analysis-based target selectors are provided in § 5.2.
5 zkFuzz: Implementation
We implemented zkFuzz for Circom in about 4,500 lines of Rust. The core implementation follows Algo. 1 with minimal heuristics. On top of this, we develop zkFuzz++, which extends the core system with static analysis-based target selectors to guide the mutation search more efficiently.
5.1 zkFuzz: Core Implementation
@@
zkFuzz begins with an enhanced parser that decomposes Circom programs into an Abstract Syntax Tree (AST) for the computation $\mathcal{P}$ and corresponding circuit constraints $\mathcal{C}$. A mutator generates program variants ($\mathcal{P}^{\prime}$), and an emulator produces execution traces for these variants using artificial inputs. A bug detector checks if these traces satisfy the original circuit constraints based on TCCT. Target selectors strategically guide input generation and program mutation towards likely bugs. Finally, mutants’ fitness scores drive selection and crossover in subsequent generations (Fig. 2).
@@
Circom Program Mutation. zkFuzz mutates Circom programs using the following methods.
@@
(b) Assignment Transformations: We randomly alter the right-hand side of assignment operators. Based on the popular mutation strategies used in mutation testing [45], we apply two types of mutations: 1) By default, we replace the right-hand expression with a random value over the finite field sampled from the skewed distribution (see § 4.2); 2) If the right-hand side contains an operator, we may also substitute it with a different operator with the same category, where we categorize the operators into arithmetic, bitwise, and logical operators.
We also implement two optional mutation methods, random value addition and expression deletion. We show detailed results in Appendix D.
Guided Search. zkFuzz guides the search by default using an error-based fitness function and a skewed distribution, as described in § 4.2.
5.2 zkFuzz++: Optimization with Target Selection via Static Analyzers
As a further optional optimization, zkFuzz++ leverages static analysis to identify promising targets, guiding both program mutation and input generation to navigate the vast search space of ZK circuits.
TABLE III: Summary of static target selectors
@@
Weak Assignment Prog, Mutation <--
White List Prog. Mutation IsZero, Num2Bits
@@
Hash-Check Input Generation h(x1) === x2
Weak Assignment Mutation. zkFuzz++ focuses on the weak assignment (<--) for the assignment transformation mutation, since <-- does not introduce constraints, and modifying its right-hand side does not affect the constraints. This optimization is based on the observation that mutating the strong assignment (<==) is ineffective, as it will likely violate the associated condition.
White List. Existing research and public audit reports have formally proven the safety of certain commonly used circuits, even if they contain weak assignments. Since mutating these circuits would immediately lead to constraint violations, zkFuzz++ allows specifying a white list of circuit templates to be skipped during the mutation phase, concentrating the mutation efforts on less well-understood or more error-prone components. In this study, we include the two most commonly used circomlib circuits with weak assignments, IsZero and Num2Bits, in the white list.
Zero-Division Pattern. As shown in § 4.2, the zero-division pattern [13] is a common source of under-constrained circuits where both the division’s numerator and denominator can be zero.
Code 6 illustrates a simple case of this issue: when both numerator and denominator are zero, the multiplication constraint (see line 7) is trivially satisfied for any value of y, due to missing checks for zero-division.
@@
(Degree 2) Consider the quadratic expression $ax^{2}+bx+c\equiv 0\bmod{q}$. A solution for $x$ is given by $x\equiv\frac{-b+\sqrt{b^{2}-4ac}}{2a}\bmod{q}$, provided that a square root of $b^{2}-4ac$ exists modulo $q$. Its existence can be verified using Euler’s criterion, and the Tonelli-Shanks algorithm [48, 49] is then employed to compute $\sqrt{b^{2}-4ac}$, yielding the solution.
@@
Hash-Check Pattern. A common pattern in ZK circuit is the hash-check pattern, where a circuit verifies that the hash of some input data matches a provided hash value (see Code 7). Naive fuzzing with program mutation may struggle to generate valid input-hash pairs, especially when cryptographic hash functions are involved. To overcome this limitation, we employ a heuristic: with a certain probability if an equality constraint (e.g., A == B) is encountered where A is an input signal that has not been assigned at this point, we update the assignment of A to match the value of B. This adjustment increases the likelihood that the input data will satisfy the hash-check constraint. A real-world example is shown in Appendix B.
@@
2 signal input x1, x2;
@@
2 signal input x1, x2;
@@
Collectively, these target selectors enable zkFuzz++ to navigate the enormous search space inherent in ZK circuits efficiently. By strategically guiding both program mutation and input generation, they significantly enhance our ability to detect critical vulnerabilities.
5.3 Support for Other DSLs
Although our current implementation primarily focuses on Circom, the core design of zkFuzz as presented in Algo 1 remains agnostic to any particular language, and we have developed an initial prototype for Noir [33], a next-generation ZK DSL gaining traction. The details of the implementation and the evaluation can be found in § 6.6.
6 Evaluation
This section presents a comprehensive evaluation designed to address the following research questions:
@@ 6 Evaluation
RQ4: What is the individual contribution of each heuristic and static analyzer in zkFuzz++ to efficient bug detection?
@@
RQ6: Is zkFuzz applicable for other DSLs like Noir?
Appendix D further considers the following questions:
@@
RQ9: What are the real-world impacts of previously unknown bugs discovered by zkFuzz?
@@
RQ10: How many over-constrained circuits occur when the constraint_assert_disabled flag is set?
@@
Benchmark Datasets. We evaluate zkFuzz on 452 real-world test suites written in Circom. Our benchmark extends the dataset from ZKAP [13] by adding new test cases from 40 additional projects, chosen based on the number of GitHub stars, forks, dependent projects, and the impact on active production services. Following [18], we classify the circuits into four categories based on the number of clauses in the constraint, denoted as $|\mathcal{C}|$. The categories are defined as follows: Small ($|\mathcal{C}|<100$); Medium ($100\leq|\mathcal{C}|<1000$); Large ($1000\leq|\mathcal{C}|<10000$); Very Large: ($10000\leq|\mathcal{C}|$).
@@
Configurations. We set a timeout of 2 hour and a maximum of 50000 generations for all our tests. We also set the number of program mutants per generation to 30, each paired with generated inputs of the same size. Program mutation and crossover operations are applied with probabilities of 0.3 and 0.5, respectively. By default, the right-hand side of a weak assignment is replaced with a random constant. However, if it consists of an operator expression, the operator is randomly substituted with a probability of 0.1. When a zero-division pattern is detected, we analytically solve for the variable and substitute it with the computed solution with a probability of 0.2. Those parameters are determined with reference to prior works on genetic algorithm mutation testing [50, 51, 52], and we empirically demonstrate the robustness of zkFuzz across different hyperparameter settings in § 6.5. We run zkFuzz with five different random seeds. All experiments were run on an Intel(R) Xeon(R) CPU @ 2.20GHz and 31GB of RAM, running Ubuntu 22.04.3 LTS.
@@
Baseline Approaches. We compare zkFuzz against four state-of-the-art Circom bug detection tools: Circomspect [20], ZKAP [13], Picus [18], and ConsCS [25]. We use two SMT solvers, Z3 [53] and CVC5 [54] for Picus. To ensure a fair comparison, we do not count warnings related to IsZero and Num2Bits as false positives, since zkFuzz explicitly whitelists these templates. Additionally, ZKAP’s unconstrained signal (US) detector flags any circuit containing unused signals as unsafe. Although the original ZKAP paper considers unused inputs problematic, our work does not regard them as bugs (see § 7). Moreover, as noted in [13], the primary source of ZKAP’s false positives is its unconstrained sub-circuit output (USCO) check, which targets unused component outputs. Therefore, we exclude both US and USCO from our main results to improve ZKAP’s precision. Note that the original Circomspect does not analyze internally called templates recursively.
@@
TABLE IV: The number of unique bugs detected by each tool, categorized by circuit size (measured by the size of constraint). TP and FP denote true positive and false positive, respectively. Prec. stands for precision, which is computed as TP / (TP + FP). Max / Min refers to the highest and lowest number of detected bugs across five different random seeds. The magnitude of TP and Precision, and FP are highlighted in blue and red, respectively. Values in the lower row of each size category, colored with gray, indicate the results only on the ZKAP dataset [13]. zkFuzz outperforms all existing methods.
Constraint Size #Bench #Bug Circomspect [20] ZKAP [13] Picus [18] (Z3 / CVC5) ConsCS [25] zkFuzz (Max / Min) zkFuzz++ (Max / Min) TP FP Prec. TP FP Prec. TP FP Prec. TP FP Prec. TP FP Prec. TP FP Prec. Small - All 181 58 49 23 0.68 36 12 0.75 22 / 23 0 1.00 14 0 1.00 57 / 57 0 1.00 58 / 58 0 1.00 Small - ZKAP 93 20 15 6 0.71 15 2 0.88 9 / 9 0 1.00 6 0 1.00 19 / 19 0 1.00 20 / 20 0 1.00 Medium - All 97 7 3 8 0.27 1 15 0.06 0 / 0 0 - 0 0 - 7 / 7 0 1.00 7 / 7 0 1.00 Medium - ZKAP 42 0 0 5 0.00 0 8 0.09 0 / 0 0 - 0 0 - 0 / 0 0 1.00 0 / 0 0 - Large - All 86 17 6 16 0.27 4 4 0.50 0 / 0 0 - 0 0 - 12 / 12 0 1.00 16 / 16 0 1.00 Large - ZKAP 50 2 2 12 0.14 0 4 0.00 0 / 0 0 - 0 0 - 0 / 0 0 1.00 1 / 1 0 1.00 Very Large - All 88 6 5 27 0.16 2 10 0.17 0 / 0 0 - 0 0 - 2 / 2 0 1.00 4 / 3 0 1.00 Very Large - ZKAP 73 5 4 15 0.21 3 0 1.00 0 / 0 0 - 0 0 - 1 / 1 0 1.00 2 / 2 0 1.00 Total - All 452 88 63 74 0.46 43 41 0.51 22 / 23 0 1.00 14 0 1.00 79 / 79 0 1.00 85 / 84 0 1.00 Total - ZKAP 258 27 21 38 0.35 18 14 0.56 9 / 9 0 1.00 6 0 1.00 20 / 20 0 1.00 23 / 23 0 1.00
6.1 RQ1: Effectiveness of zkFuzz